OpenStudy (anonymous):

solve this using chain rule y= square root 1+ square root x show ur solutions plss

zepdrix (zepdrix):

\[\large y=\sqrt{1+\sqrt x}\]This is the problem?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Change your roots to fractional exponents. Then work from the outside inward.\[\large y=(1+x^{1/2})^{1/2}\]The outermost function is the ( ) ^ 1/2 We'll differentiate that first, then apply the chain rule.\[\large y'=\frac{1}{2}(1+x^{1/2})^{-1/2}\cdot (1+x^{1/2})'\]The prime is to show that we still need to differentiate that part. That copy of the inside came about due to the chain rule.

OpenStudy (anonymous):

ahh i know it

zepdrix (zepdrix):

k cool c:

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

how about y=9/x^3 + 3/2x^2 find y^2 using higher derivatives